OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form. 3x – 3y = –8 x – 3y = –9 x – 3y = –8 3x + y = 9

OpenStudy (anonymous):

I got this answer 3x+y= 3 ? but it does not match with what I have in the answers

OpenStudy (cwrw238):

slope = 4-3 / 4-1 = 1/3 y-y1 = 1/3 (x - x1) x1,y1 = 1,3 y - 3 = 1/3(x - 1) 3y - 9 = x - 1 3y - x = 8 or x - 3y = -8

OpenStudy (anonymous):

I did some stps super wrong thank u!